THEORETICAL STUDIES OF THE ELECTRONIC STRUCTURE OF THE GROUND AND LOW-LYING EXCITED STATES OF ScH_{2}^{++}, TiH_{2}^{++}, VH_{2}^{++}, CrH_{2}^{++}, and MnH_{2}^{++}

\underline{JAMES F. HARRISON}, JOSEPH KENNY, Department of Chemistry and Center for Fundamental Materials Research, Michigan State University, East Lansing MI 48824-1322.

In all of the titled molecules the two positive charges remain localized essentially on the transition metal and the low-lying electronic states are those of the corresponding dipositive free ion split by a virtually intact, polarized H_{2} molecule. We discus the character of these states, their binding energy, and the splitting of the free ion term as determined by high level ab-initio calculations.